OpenStudy (anonymous):

Without using a calculator, find the exact value as fraction (not a decimal approximation). 5Pi/4

OpenStudy (anonymous):

Multiply it by 180/pi

OpenStudy (anonymous):

AMBROSE I NEED HELP QUICK

OpenStudy (anonymous):

pis cancel you get 5(180)/4 = 900/4 = 225